OpenStudy (anonymous):

Find the limit for the given function. lim θ→0 sin(cos 4θ) sec 5θ

OpenStudy (anonymous):

\[\lim_{\theta\to0}\sin(\cos4\theta)\sec5\theta=\lim_{\theta\to0}\frac{\sin(\cos4\theta)}{\cos5\theta}=\frac{\sin(\cos0)}{\cos0}\]